OpenStudy (xhawaiiangirlx):

When solving negative one over three.(x − 15) = −4, what is the correct sequence of operations? Multiply each side by -1/3 , subtract 15 from each side Multiply each side by −3, add 15 to each side Multiply each side by -1/3, add 15 to each side Multiply each side by 3, subtract 15 from each side @pooja195

OpenStudy (edisonvalentimlopes):

Is this the equation? \[-\frac{ 1 }{ 3 }\times (x-15) = -4\]

OpenStudy (xhawaiiangirlx):

yes

OpenStudy (xhawaiiangirlx):

but do you know the answer?

OpenStudy (edisonvalentimlopes):

Yes.... but directly? Yes..... B, but do you know WHY?

OpenStudy (xhawaiiangirlx):

im not sure

OpenStudy (edisonvalentimlopes):

Our goal is x = ...., ok? First, multiply each side by -3: \[(-3)\quad -\frac{ 1 }{ 3 }\times (x-15)=-4 \to (x-15)=12\]

OpenStudy (edisonvalentimlopes):

ok?

OpenStudy (edisonvalentimlopes):

Second, add 15 to each side:\[(+15)\quad x-15=12 \to x=27\]

OpenStudy (edisonvalentimlopes):

Do you understand?

OpenStudy (xhawaiiangirlx):

yes

OpenStudy (xhawaiiangirlx):

thank you i got 100% thanks to you!!

OpenStudy (edisonvalentimlopes):

Thanks for your trust and, If you need, just send me a message or an alert.